How do i use the substitution method for
-3x+y=4
x+13y=22

Answer by algebrahouse.com(1659)

How do i use the substitution method for
-3x + y = 4------->y = 3x + 4 {added 3x to both sides}
x + 13y = 22

x + 13(3x + 4) = 22 {substituted 3x + 4, in for y, into bottom equation}
x + 39x + 52 = 22 {used distributive property}
40x + 52 = 22 {combined like terms}
40x = -30 {subtracted 52 from both sides}
x = -3/4 {divided both sides by 40}

y = 3x + 4
y = 3(-3/4) + 4 {substituted -3/4, in for x, into y = 3x + 4}
y = -9/4 + 4 {multiplied 3 by -3/4}
y = -9/4 + 16/4 {got common denominator}
y = 7/4 {combined like terms}
y = 1 3/4 {wrote as a mixed number}

x = -3/4 and y = 1 3/4
www.algebrahouse.com